### v1

@@ Abstract
Purpose Computational head models are essential tools for studying the risk of mild traumatic brain injury (mTBI) under different activities and across populations. However, different computational models incorporate varied levels of anatomical details, such as cortical folds. In this study, we aim to determine the effect of modeling cortical folds on mTBI risk assessment.
Methods We compared the gyrencephalic (with cortical folds) and lissencephalic (without cortical folds) FE models of 18 subjects aged 9 - 18 years, under a rotational head acceleration event. A rotational acceleration of 10 krad/s2 and 10 ms duration was simulated about each principal head axis. We analyzed different mTBI injury metrics, including maximum principal strain (MPS95), maximum principal strain rate (MPSR95), and cumulative strain damage measure (CSDM15), for the whole brain as well as for specific regions of interest (ROIs).
Results Modeling cortical folds consistently predicted higher injury metrics across all individuals and rotational direction, with the bias (mean $\pm$ std. dev.) of $15.1\pm 6.5\%$ in MPS95, $12.9\pm 5.6\%$ in MPSR95, and $8.8\pm 11.09\%$ in CSDM15. We also find that the regions of high strain concentrations vary significantly between the two models, with the DICE metric on peak MPS ranging between 0.070.430.07-0.43 and DICE on CSDM15 ranging between 0.420.700.42-0.70. Modeling cortical folds also affects injury metrics in all ROIs, even the ones that remain geometrically unaltered in the two model types, such as the corpus callosum, cerebellum, and brain stem.
Conclusions The study finds that modeling cortical folds significantly alters the region of high brain deformations and the mTBI risk under head rotations.
@@ 1 Introduction
Mild traumatic brain injury (mTBI), or concussion, is a highly prevalent and heterogeneous condition, affecting different individuals differently [1]. Computational head models have been used to study the mTBI risk of various activities in different populations [2, 3, 4, 5, 6, 7]. These head models vary significantly in terms of capturing the level of anatomical details, the individual variations in these details, and the choice of finite element (FE) mesh type or material models. Cortical folds, also known as the gyri and sulci in the brain’s cerebrum, are one such anatomical detail. However, the effect of incorporating cortical folds into FE head models on mTBI risk assessment is not well understood. In this paper, we aim to study the effects of modeling subject-specific cortical folds on the mTBI risks under head rotational accelerations.
@@
Previous studies have investigated the role of cortical folds on mTBI injury metrics, such as strains [8, 9, 10, 11, 12] and stresses [10, 12, 8] in the brain tissues. These computational [8, 9, 10, 12] and experimental [11] studies compared the brain strains and stresses developed in a gyrencephalic (with cortical folds) model and a lissencephalic (smooth cortical surface without folds) model. The findings of these studies show conflicting trends regarding which model predicted higher strains or stresses. Some computational studies on 3D human head models [10, 9] and 2D-sagittal slice models [8] found higher strain and stress in the lissencephalic models. On the other hand, an experimental study on 2D coronal surrogates [11] and a computational study on 2D-axial model [12] found higher strains in gyrencephalic models. These discrepancies could stem from differences in the loading conditions, tissue material properties, and the variations in the cortical folds modeled across the studies. In this study, we investigate the role of cortical folds on mTBI response, while accounting for cortical fold variations across different individuals and the directions of loading.
@@
Several mesoscale studies at the level of gyri have explored the effects of cortical fold parameters on the stress and strain distributions [13, 14, 15, 16]. These mesoscale studies are based on 2D plane-strain computational models of simplified gyri and sulci geometry representing brain cortical folds. These studies investigated the impacts of sulcus depth, width, and thickness on the stress and strain distributions. While most of the studies found that the strain increases with the sulcal depth [15, 13, 14], one study found higher strains in lissencephalic models [16]. However, since the variations in the cortical folds were not based on actual human brain anatomy, the practical and statistical significance of incorporating cortical folds in subject-specific FE head models for mTBI risk assessment remains unanswered, and the computational studies on mTBI continue to use both lissencephalic and gyrencephalic models to understand mTBI risk in practical scenarios [2, 3].
In this paper, we study the effect of modeling subject-specific cortical folds on mTBI risk by comparing subject-specific gyrencephalic and lissencephalic models of 18 subjects under rotational head motion. We study the differences in several mTBI metrics for the two model types for all individuals. The results of this study provide the importance of incorporating cortical fold details in large-population studies, where modeling cortical folds might present significantly higher model development efforts.
@@
The paper is organized as follows. Section 2 describes the methods and workflows used in developing the subject-specific models (gyrencephalic and lissencephalic) from their medical images, and the details of the finite element simulations. Section 3 provides the results on different mTBI metrics from the computational study, and Section 4 discusses the implications of these results on our understanding and future steps.
@@ 2 Materials and Methods
In this section, we describe the development of gyrencephalic (with cortical folds) and lissencephalic (without cortical folds) finite element (FE) head models (Section 2.1), followed by the details of FE simulation (Section 2.2), injury metrics (Section 2.3), and data analyses (Section 2.4) to understand the differences between the mTBI risk assessment using the two model types.
@@ 2.1 Subject-specific Head Modeling
We developed the FE head models of 18 individuals (8 males, 10 females; 9–18 years old) from their magnetic resonance imaging (MRI) scans. The number of subjects by age and sex in each group is provided in Table 1. In this section, we describe the steps for developing the subject-specific gyrencephalic and lissencephalic models from the medical images.
@@
Table 1: Number of subjects by age group and gender
@@ 2.1.1 Medical Image Segmentation
The first step in generating a medical image-based FE model is medical image segmentation, which delineates the boundary of different parts of the model. The medical images of each individual consisted of structural MRI (T1-w) and diffusion tensor imaging (DTI) (Fig. 1a). The T1-structural MRIs were segmented into the skull, the meninges (dura, falx, and tentorium), sub-arachnoid cerebral spinal fluid (CSF), ventricles, and different regions of the brain. The segmentation of the brain tissue into the cortical gray matter, white matter, brain stem, and deep brain regions was obtained using the FreeSurfer software package [17]. The resulting gray matter incorporates cortical folds, or gyri and sulci, in 68 parcellations (Fig. 1b), providing a gyrencephalic head model segmentation. The detailed steps of the MRI pre-processing and segmentation workflow is provided in A.
@@
The lissencephalic (without cortical folds) models were generated by modifying the detailed gyrencephalic model segmentations. First, the heterogeneous sub-arachnoid CSF in the gyrencephalic segmentation was replaced by a uniform 1 mm thick layer of CSF lining the meninges. The cortical gray and white matter matter with gyri and sulci in the gyrencephalic segmentation were replaced by uniform 4 mm thick gray matter and the remaining space between the gray matter and the deep brain structures filled with new white matter, to provide the lissencephalic model segmentations (Fig. 1b).
@@ 2.1.2 Finite Element Meshing
The FE mesh was generated directly from the segmentation by converting the segmented image voxels (1 mm $\times$ 1 mm $\times$ 1 mm) directly to hexahedral elements using a custom MATLAB script (Mathworks Inc.) (Fig. 1c). The voxel meshing allowed fast mesh generation (<15<15minutes on 1 CPU) while retaining fine subject-specific details, such as cortical folds. On the other hand, the interfaces between two materials were jagged and non-smooth. However, studies have found these numerical approximations not to cause significant issues [18, 19, 20, 21]. A reduced integration scheme with hourglass control was implemented to prevent volumetric locking behavior when modeling nearly incompressible materials.
@@ 2.1.3 Material Modeling
The brain tissue was modeled as a hyperelastic material with the axonal tract orientation from the DTI scans incorporated through an anisotropic hyperelastic material model. The time dependence is captured using a first-order Prony series. The CSF was modeled using the Mie Gruneisen equation of state, and time dependence was modeled as Newtonian viscous flow. The meninges were modeled as linear elastic solid, and skull as a rigid solid. The material model definitions of the brain tissue, CSF, and the meninges were based on a previously validated FE head models [5, 22]. The experimental evaluation of one of the subject-specific FE models used in this study is presented in the Supplementary Sections S2 and S3.
@@ 2.2 Finite Element Simulations
Each finite element head mode was used to simulate three head acceleration events, where a half-sinusoidal rotational acceleration pulse was applied to the skull about a principal anatomical axis (axial, sagittal, and coronal) [23]. The peak acceleration was selected to be 10 krad/s2 for a duration of 10 ms, resulting in a peak angular velocity of 60 rad/sec (Fig 1d), to simulate a concussive event [24]. The simulations were conducted using an explicit time integration scheme in Abaqus FE software (Simulia, Dassault), and the results from these simulations were compared to study the effect of cortical folds on mTBI risk assessment.
Figure 1: Workflow for studying the effect of modeling cortical folds on mTBI injury metrics: a) subject specific magnetic resonance imaging (MRI) scans, b) Medical image processing to create gyrencephalic and lissencephalic head segmentations, c) Voxel meshing, d) Idealized head rotation acceleration about three cardinal planes in separate simulations, e) Post-processing to obtain injury metrics
@@ 2.3 Injury Metrics
To quantify the effect of cortical folds on mTBI risk assessment, we compare different injury metrics obtained from the gyrencephalic and lissencephalic model simulations for each subject.
The FE simulations in Abaqus/Explicit provided the maximum principal logarithmic strain (MPS) and maximum principal logarithmic strain rate (MPSR) at each material (element) and time point [23]. An element was flagged as damaged after the MPS at its location exceeded 0.15 [23]. A cumulative strain damage measure (CSDM15) was calculated as the fraction of the damaged elements’ volume to the total brain tissue volume. Similarly, we also calculated the peak maximum principal strain rate (PMPSR) for all the elements in a model, and cumulative strain damage measure (CSRDM50) based on elements exceeding PMPSR of 50 s-1.
To study the effect of cortical folds on the overall global response of the brain during the head acceleration event, we compare the highest strains and strain rates developed anywhere within the brain tissue at any point during the simulation in the gyrencephalic and lissencephalic models of a given subject. The FE simulations were post-processed to obtain the 95th percentile of the MPS (MPS95) and MPSR (MPSR95) to provide the highest MPS and MPSR within the whole brain (instead of the 100th percentile, which is commonly avoided in complex biological FE models to avoid numerical artifacts).
Since the peak MPS (or MPSR) at a given anatomical point in the brain occurs at different times in the two models, we calculated the peak MPS (PMPS) experienced by the element over the whole duration of the head acceleration simulation (30 ms) [25]. PMPS allowed the comparison of the peak strain experienced by the elements at the same anatomical locations in the two models, i.e. the similarity of the PMPS distribution.
@@ 2.4 Data Analysis
We use Sørensen DICE coefficient (DICE) to compare the similarity of the spatial distributions in the gyrencephalic and lissencephalic models. The DICE coefficient was obtained for PMPS and CSDM, as defined below:
$\text{DICE}=\frac{2|G\cap L|}{|G|+|L|}$ (1)
where $G$ and $L$ are the element sets in the gyrencephalic and the lissencephalic models, $|G\cap L|$ is the intersection of the two sets, i.e. elements at corresponding spatial locations in both models that have the same PMPS (or CSDM) values. We consider an element to have the same values in the two models if the absolute difference is less than 0.025. A higher value of DICE indicates a better spatial agreement between the two models, with 1 being all elements have the same values and 0 being no elements have the same values.
The DICE coefficient for PMPS represents the similarity over the whole range of strain values, whereas the DICE coefficient for the CSDM represents the similarity in damaged elements in the two models.
We compare the cohort-level means and standard deviations of the peak MPS95, MPSR95, and CSDM from the gyrencephalic to the lissencephalic models. We perform this comparison across the whole brain parenchyma, and specific anatomical regions of interest (ROI), including cerebellum, brain, corpus callosum, cerebral gray matter, cerebral white matter, and deep brain region comprising the thalamus, hippocampus, amygdala, putamen, pallidum, and hypothalamus. These comparisons capture the peak values of the injury metrics, without accounting for the precise spatial locations of the maximum strain within each ROI.
We performed the Bland-Altman test to obtain the bias between the gyrencephalic and lissencephalic models for peak MPS95, peak MPSR95, and CSDM15. We normalize the bias by the maximum gyrencephalic result for the case to understand the practical significance of the bias. We performed a statistical significance test using paired t-test and $p<0.005$ is considered significant.

### v2

@@ Abstract
Purpose Computational head models are essential tools for predicting the risk of mild traumatic brain injury (mTBI) under different activities and across populations. However, different computational models incorporate varied levels of anatomical details, such as cortical folds. In this study, we aim to determine the effect of modeling cortical folds on mTBI risk assessment.
Methods We compared gyrencephalic (with cortical folds) and lissencephalic (without cortical folds) FE models of 18 subjects aged 9 - 18 years under a rotational head acceleration event. A rotational acceleration of 10 krad/s2 and 10 ms duration was simulated about each principal head axis. We analyzed different tissue-level mTBI injury metrics, including maximum principal strain (MPS95), maximum principal strain rate (MPSR95), and cumulative strain damage measure (CSDM15), for the whole brain and for specific regions of interest (ROIs).
Results The inclusion of cortical folds consistently yielded higher injury metrics across all individuals and rotational directions, with a bias (mean $\pm$ std. dev. relative to maximum lissencephalic values) of $21.7\pm 9.1\%$ in MPS95, $17.1\pm 7.6\%$ in MPSR95, and $14.4\pm 11.3\%$ in CSDM15. Modeling cortical folds significantly affected the spatial strain distributions, with the DICE similarity coefficient on peak MPS ranging between 0.070.430.07-0.43 and on CSDM15 ranging between 0.420.700.42-0.70; and increasing the peak injury metrics even in the geometrically unaltered regions of interest, such as the corpus callosum, cerebellum, and brain stem, by up to $\sim$50%. Conclusions The study finds that the inclusion of cortical folds significantly alters the pattern of deformation in the brain, thereby affecting the mTBI risk predictions head rotations.
@@ 1 Introduction
Mild traumatic brain injury (mTBI), including concussion, is a highly prevalent and heterogeneous condition, affecting different individuals differently [1]. Computational head models have been used to study the mTBI risk of various activities in different populations [2, 3, 4, 5, 6, 7]. These head models vary significantly in terms of the incorporated level of anatomical detail, the individual anatomical variations, and the choice of finite element (FE) mesh type or material models. Cortical folds, which consist of the gyri and sulci in the brain’s cerebrum, are one such anatomical detail. While some studies have used models with cortical folds to study the risk of mTBI [8, 7, 9, 3], others use FE models without cortical folds [10, 6, 2]. However, the effect of incorporating cortical folds into FE head models on mTBI risk assessment is not well understood. In this paper, we aim to study the effects of modeling several subject-specific cortical folds on predicting the mTBI risk under head rotational accelerations.
@@
Previous studies have investigated the role of cortical folds on tissue level mTBI injury metrics, such as strains [11, 12, 13, 14, 15] and stresses [13, 15, 11] in the brain tissue. These computational [11, 12, 13, 15] and experimental [14] studies compared the brain strains and stresses developed between a gyrencephalic (with cortical folds) and a lissencephalic (smooth cortical surface without folds) model, but resulted in conflicting findings. Some computational studies based on 3D [12] and 2D sagittal models [11] of the human head, and 3D head models of multiple species [13] reported higher strain and stress in the lissencephalic models. On the other hand, an experimental study on 2D coronal surrogates of porcine brain [14] and a computational study on 2D-axial human brain model [15] found higher strains in gyrencephalic models. These discrepancies could stem from differences in the impact loading conditions, tissue material properties, and the geometrical variations in the cortical folds modeled across the studies.
@@
Some computational studies have explored the role of geometrical variations in cortical folds on the deformation in the brain tissue [16, 17, 18, 19]. These studies are based on 2D plane-strain mesoscale models of a square region incorporating approximately a couple of gyri with simplified geometry. These studies also reported conflicting trends, with most of the studies finding higher strains in gyrencephalic models [17, 16, 19], but one study reporting higher strains in the lissencephalic model [18]. The models used in these studies varied in terms of the applied boundary conditions to approximate an impact loading, with some studies applying uniform compression on different sides of the mesoscale model [19, 18, 17], and others applying non-uniform shear acceleration field obtained from a full head simulation [16]. The studies also had different tissue mechanical properties. Also, since the variations in the cortical folds were not based on actual human brain anatomy, the practical and statistical significance of incorporating cortical folds in subject-specific FE head models for mTBI risk assessment remains unanswered, and the computational studies on mTBI continue to use both lissencephalic and gyrencephalic models to understand mTBI risk in practical scenarios [2, 3].
In this study, we investigate the role of cortical folds on mTBI risk, while accounting for cortical fold variations across different individuals and the directions of loading. We develop subject-specific gyrencephalic and lissencephalic models of 18 subjects and compare their mTBI response under head rotations about the three principal anatomical axes.The results of this study will provide the importance of incorporating cortical fold details in FE head models used for mTBI risk prediction.
@@
The paper is organized as follows. Section 2 describes the methods and workflows used in developing the subject-specific models (gyrencephalic and lissencephalic) from their medical images and the details of the finite element simulations. Section 3 provides the results on different mTBI metrics from the computational study, and Section 4 discusses the implications of these results on our understanding and future steps.
@@ 2 Materials and Methods
In this section, we describe the development of gyrencephalic (with cortical folds) and lissencephalic (without cortical folds) finite element (FE) head models (Section 2.1), followed by the details of FE simulation (Section 2.2), injury metrics (Section 2.3), and data analyses (Section 2.4) to detail the differences between the mTBI risk assessment using the two model types.
@@ 2.1 Subject-specific Head Modeling
We developed the FE head models of 18 individuals (8 males, 10 females; 9–18 years old) from their magnetic resonance imaging (MRI) scans. The number of subjects by age and sex in each group is provided in Table 1. Subject-specific gyrencephalic and lissencephalic models were generated directly from the medical images.
@@
Table 1: Number of subjects by age and sex
@@ 2.1.1 Medical Image Segmentation
The first step in generating a medical image-based FE model is medical image segmentation, which delineates the boundary of different parts of the model. The medical images of each individual consisted of structural MRI (T1-w) and diffusion tensor imaging (DTI) (Fig. 1a).
The T1-structural MRIs were first pre-processed to correct for Gibbs-ringing and bias-field artifact using open source software MRtrix [20] and ANTS [21], respectively. All MRI scans were registered rigidly to the MNI152 atlas using 3D slicer (http://www.slicer.org) [22]. The segmentation of the brain tissue into the cortical gray matter, white matter, brain stem, and deep brain regions was obtained using the FreeSurfer software package [23]. The resulting gray matter incorporates cortical folds, or gyri and sulci, in 68 parcellations (Fig. 1b), providing a gyrencephalic head model segmentation. Automated brain segmentation on the enhanced T1 MRIs using the FSL software [24] provided the skull segment. The additional 1 mm thick meninges (dura, falx, and tentorium) segment was added manually after combining the FreeSurfer and FSL segmentations using 3D Slicer. The sub-arachnoid cerebrospinal fluid (CSF) was filled in the space between the meninges and the brain segments. A one mm layer of CSF was added around the meninges to ensured that brain tissue was surrounded by the CSF everywhere. This workflow provided the gyrencephalic (with cortical folds) model segmentations.
@@
The lissencephalic models were generated by modifying the detailed gyrencephalic model segmentations. First, the heterogeneous sub-arachnoid CSF in the gyrencephalic segmentation was replaced by a uniform 1 mm thick layer of CSF lining the inner surface of the meninges. The cortical gray and white matter matter with gyri and sulci in the gyrencephalic segmentation were replaced by a uniform 4 mm thick gray matter layer inside the new uniform CSF layer segment. The remaining space between the gray matter and the deep brain structures was filled with white matter, to provide the lissencephalic model segmentations (Fig. 1b).
@@ 2.1.2 Finite Element Meshing
The FE mesh was generated directly from the segmentation by converting the segmented image voxels (1 mm $\times$ 1 mm $\times$ 1 mm) directly to hexahedral elements using a custom MATLAB script (Mathworks Inc.) (Fig. 1c). The voxel meshing allowed fast mesh generation ($\sim 15$ minutes on 1 CPU) while retaining fine subject-specific details, such as cortical folds. While the interfaces between two materials were jagged and non-smooth, a recent study found that the voxel and conformal mesh (with smooth interfaces) result in identical response when comparing strain based injury metrics, such as the 99th percentile strain or the volume fraction of brain with strains over certain thresholds [25]. A number of FE head models that incorporate cortical folds use voxel mesh [26, 27, 28]. A reduced integration scheme with hourglass control was implemented (element C3D8R) to prevent volumetric locking behavior given the nearly incompressible material behavior of the brain tissues.
@@ 2.1.3 Material Modeling
The brain tissue was modeled as an anisotropic visco-hyperelastic material with the axonal tract orientation from the DTI scans. The time dependence is captured using a first-order Prony series. The bulk response of the CSF was modeled using the Mie Gruneisen equation of state, and the shear response as Newtonian viscous flow. The meninges were modeled as a linear elastic solid, and skull as a rigid solid. The material model definitions of the brain tissue, CSF, and the meninges were based on a previously published FE head model [5, 29]. The evaluation of the brain mechanical response one of the subject-specific FE models used in this study against experimental data is presented in the Supplementary Sections S2 and S3.
@@ 2.2 Finite Element Simulations
Each finite element head model was used to simulate three head acceleration events, where a half-sinusoidal rotational acceleration pulse was applied to the skull about a principal anatomical axis (axial, sagittal, and coronal) [30]. The peak acceleration was selected to be 10 krad/s2 for a duration of 10 ms, resulting in a peak angular velocity of 60 rad/sec (Fig 1d). The loading magnitude was selected based on measured data in sports such as mixed martial arts and American football, to represent an event with a high probability of concussion [31, 32]. The simulations were conducted using an explicit time integration scheme in Abaqus FE software (Simulia, Dassault).
Figure 1: Workflow for studying the effect of modeling cortical folds on mTBI injury metrics: a) Subject-specific magnetic resonance imaging (MRI) scans, b) Medical image segmentation to create gyrencephalic and lissencephalic head models, c) Voxel meshing, d) Idealized head rotation acceleration profile applied about the three cardinal anatomical planes, and e) Post-processing to obtain injury metrics
@@ 2.3 Injury Metrics
To quantify the effect of cortical folds on mTBI risk assessment, we compare different tissue-level injury metrics obtained from the gyrencephalic and lissencephalic model simulations for each subject. Previous studies have used the strain and strain rate in the brain tissue from FE simulations with mTBI risk [33, 34, 35, 36]. The FE simulations in Abaqus/Explicit provided the maximum principal logarithmic strain (MPS) and maximum principal logarithmic strain rate (MPSR) at each material (element) and time point [30].
An element was flagged as damaged after the MPS at its location exceeded 0.15, based on experimental neuronal injury thresholds [37, 38, 39, 40, 41, 42]. A cumulative strain damage measure (CSDM15) was calculated as the fraction of the damaged elements’ volume to the total brain tissue volume. We also calculated the cumulative strain rate damage measure (CSRDM50) based on elements exceeding MPSR of 50 s-1 for a representative case.
Since the peak MPS at a given material point occurs at different times in the two models, we calculated the time-invariant peak strain distribution for a given simulation as the peak MPS experienced by an element over the whole duration of the head acceleration simulation (30 ms) [43]. We refer to this time-invariant peak MPS distribution as just ‘the MPS distribution’ in the following sections, unless otherwise stated. This allows the comparison of peak strain experienced at the same anatomical locations between the two models, i.e. the similarity of strain distribution. We also calculated the time-invariant peak MPSR distributions for a representative case.
To study the effect of cortical folds on the overall global response of the brain during a head acceleration event, we compare the highest strains and strain rates developed anywhere within the brain tissue at any point during the simulation in the gyrencephalic and lissencephalic models of a given subject. The FE simulations were post-processed to provide the 95th percentile of the MPS (MPS95) and MPSR (MPSR95) (instead of the 100th percentile, which is commonly avoided in complex biological FE models to avoid numerical artifacts).
@@ 2.4 Data Analysis
We use Sørensen DICE coefficient (DICE) to compare the similarity of the spatial distributions in the gyrencephalic and lissencephalic models. The DICE coefficient was obtained for the MPS and CSDM distributions, as defined below:
$\text{DICE}=\frac{|G\cap L|}{|G|}$ (1)
where $G$ and $L$ are the element sets in the gyrencephalic and the lissencephalic models, $|G\cap L|$ is the intersection of the two sets, i.e. elements at corresponding spatial locations in both models that have the same MPS (or CSDM) values. We consider an element to have the same values in the two models if the absolute difference is less than 0.025. The DICE value is sensitive to the threshold, and we selected the value of 0.025 to be consistent with the contour plots showing MPS distribution in the results section (Section 3). A higher value of DICE indicates a better spatial agreement between the two models, with 1 indicating all elements have the same values and 0 being no elements have the same values. The DICE coefficient for MPS represents the similarity over the whole range of strain values, whereas the DICE coefficient for the CSDM represents the similarity in damaged elements in the two models.
We calculate the means and standard deviations of the peak values of the injury metrics (MPS95, MPSR95, and CSDM15) from all subjects to compare the gyrencephalic and lissencephalic models. We perform this comparison across the whole brain, and specific anatomical regions of interest (ROI), including cerebellum, brain, corpus callosum, cerebral gray matter, cerebral white matter, and deep brain region comprising the thalamus, hippocampus, amygdala, putamen, pallidum, and hypothalamus. We performed the Bland-Altman test to obtain the bias between the gyrencephalic and lissencephalic models for peak MPS95, peak MPSR95, and CSDM15. We normalize the bias by the maximum lissencephalic result for the case to understand the practical significance of the bias. We performed a statistical significance test using paired t-test, and $p<0.005$ is considered significant.